73. CH 3 OH can be synthesized by the reaction: CO(g)+2 H 2 (g) CH 3 OH(g) What volume of H 2 gas (in ), at 748 mmHg and 86 degrees * C , is re quired to synthesize 25.8 g CH 3 OH ? How many liters of CO gas, measured under the same conditions , are required?
Answer #1

Given reaction

CO(g) + 2 H​​​​​​2 (g) ----------------> CH3OH (g)

2*2 g of Hydrogen on reaction with excess of CO produced 32 g of CH3OH

Weight of Hydrogen required to get 25.8 g of CH3OH

= 25.8 g * 2* 2 g/32 g = 3.225 g

Number of moles of Hydrogen gas = Weight of Hydrogen/ Molar mass of Hydrogen

= 3.225 g/( 2 g/mol) = 1.6125 moles

Given conditions are

Temperature = 86°C = 86 +273 = 359 K

Pressure = 748 mm Hg = 0.984 atm

Gas constant (R) = 0.0821 L-atm/(mol-K)

V = nRT/P

= 1.613 moles * 0.0821 L-atm/(mol-K) * 359 K/0.984 atm

= 48.3 L

32 g of CH3OH is produced by the reaction of 28 g of CO

25.8 g of CH3OH is produced by the reaction of how many grams of CO

= 25.8 g *28 g) /32 g = 22.575 g

Number of moles of CO is 22.575 g/ 28 g/mol <

= 0.81 moles

Given conditions are

Temperature = 86°C = 86 +273 = 359 K

Pressure = 748 mm Hg = 0.984 atm

Gas constant (R) = 0.0821 L-atm/(mol-K)

V = nRT/P

= 0.81 Moles * 0.0821 L-atm/(mol-K) * 359 K/0.984 atm

= 24.26 L